OpenStudy (anonymous):

can someone help me with 9 a,d,g and 10 B,E,F

OpenStudy (anonymous):

can someone help me with 9 a,d,g and 10 b,e,f

OpenStudy (kymber):

I can totally read that

OpenStudy (anonymous):

the imagin not clear

OpenStudy (kymber):

Y u no use scanner.

OpenStudy (anonymous):

can uread direction at least? i can write out what i need

OpenStudy (kymber):

Just type it, buddy.

OpenStudy (anonymous):

okay 9a ) f(x)=x(7-x) d) f(x)=-(x-3)(x+5) g) f(x)=2(x-3)(x-8)

OpenStudy (anonymous):

find coordinates of x/ y intercept and max or mini point on the grapg of these quad fuction. show answer by obtaining reasoing with sumbolic forms

OpenStudy (kymber):

I suggest putting it into a new question